Determine the molality of water at 25 degree centigrade.(explain)

Grade:12

1 Answers

Gaurav
9 years ago
Lets find out the molarity of 1 L sample of water
Volume of water = 1L
Density of water = 1 kg/L
Mass of 1 L water = 1 kg = 1000g
Number of moles of water = mass of water / molar mass of water = 1000/18
Molarity = number of moles/ volume = 1000/18/1 = 55.55 m